Assume that blood pressure readings are normally distributed with a mean of 117 and a standard deviation of 5. A researcher wishes to select people for a study but wants to exclude the top 10 percent. What would be the cutoff blood pressure to qualify for the study?
a)120.5
b)117.0
c)124.6
d)123.4
e)121.8

Answer by stanbon(75887)

Assume that blood pressure readings are normally distributed with a mean of 117 and a standard deviation of 5. A researcher wishes to select people for a study but wants to exclude the top 10 percent. What would be the cutoff blood pressure to qualify for the study?
Find the z-value with a left-tail of 0.90
invNorm(0.90) = 1.6449
---
Find the corresponding bp value:
bp = z*s + u
---
blood pressure = 1.6449*5+117 = 125.22
===============
Cheers,
Stan H.
===============




a)120.5
b)117.0
c)124.6
d)123.4
e)121.8